a woman jogging at 6 miles an hour passes a man riding his bicycle. What does this mean? my e-mail address is laci90h@yahoo.com
Answer by Alan3354(69443)

a woman jogging at 6 miles an hour passes a man biking in the opposite direction at 12 miles an hour. if they maintain their speeds, how far from each other will they be 10 minutes after passing?
-----------------
Their speeds are added when going in opposite directions. 6+12 = 18 mph.
10 min = 1/6 hour
18*(1/6) = 3 miles